# NLP for Social Good: A Survey of Challenges, Opportunities, and Responsible Deployment

### v1

@@ Abstract
Recent advancements in large language models (LLMs) have unlocked unprecedented possibilities across a range of applications. However, as a community, we believe that the field of Natural Language Processing (NLP) has a growing need to approach deployment with greater intentionality and responsibility. In alignment with the broader vision of AI for Social Good Tomašev et al. (2020), this paper examines the role of NLP in addressing pressing societal challenges. Through a cross-disciplinary analysis of social goals and emerging risks, we highlight promising research directions and outline challenges that must be addressed to ensure responsible and equitable progress in NLP4SG research.
NLP for Social Good: A Survey of
@@
Antonia Karamolegkou1, Angana Borah2, Eunjung Cho3, Sagnik Ray Choudhury4, Martina Galletti5,6, Rajarshi Ghosh7, Ishani Mondal8, Pranav Gupta9, Oana Ignat10, Priyanka Kargupta11, Neema Kotonya12, Hemank Lamba12, Sun-Joo Lee13, Arushi Mangla10, Deniz Nazarova14, Poli Nemkova4,13, Dina Pisarevskaya15, Naquee Rizwan16, Nazanin Sabri17, Dominik Stammbach18, Anna Steinberg19,20, David Tomás21, Steven R Wilson22, Bowen Yi2,23, Jessica Zhu8, Arkaitz Zubiaga15, Anders Søgaard1, Alexander Fraser20,27, Zhijing Jin3,24,25,26, Rada Mihalcea2, Joel R. Tetreault12, Daryna Dementieva27 1University of Copenhagen 2University of Michigan-Ann Arbor 3ETH Zurich 4 University of North Texas 5Sony Computer Science Laboratories - Paris 6University of Rome “La Sapienza” 7Algoverse 8University of Maryland, College Park 9Lowe’s 10Santa Clara University 11University of Illinois Urbana-Champaign 12Dataminr 13 United Nations Development Programme (UNDP) 14University of Washington 15Queen Mary University of London 16IIT Kharagpur 17University of California San Diego 18 Princeton University 19LMU Munich 20Munich Center for Machine Learning (MCML) 21University of Alicante 22University of Michigan-Flint 23University of Southern California 24Max Planck Institute for Intelligent Systems, Tübingen 25Vector Institute 26University of Toronto 27Technical University of Munich Correspondence: antka@di.ku.dk, daryna.dementieva@tum.de
@@ 1 Introduction
To fully realize the potential of NLP, it is essential to look beyond technical achievements and reframe tasks around pressing societal needs. We draw on insights from the United Nations Sustainable Development Goals [^1] https://sdgs.un.org/goals (UN SDGs) and the 2025 Global Economic Risks Report [^2] https://www.weforum.org/publications/global-risks-report-2025/digest (GR) to provide a foundation for an interdisciplinary recontextualization of NLP, encouraging reflection on how language technologies intersect with today’s most pressing challenges. We selected these two agendas as, from a social good perspective, UN SDGs offer a global framework for fostering peace and prosperity for people and the planet. However, while highly influential, these goals were established in 2015—prior to the rapid advancements in artificial intelligence. To contextualize them within today’s technological landscape, we also draw on insights from the 2025 GR Report, which highlights both the transformative potential and the emerging global risks (GRs) associated with technology and information processing. Our effort builds on prior research that assesses the role of NLP through positive impact Hovy and Spruit (2016); Jin et al. (2021), maps NLP4SG work to the SDGs Adauto et al. (2023); Gosselink et al. (2024), outlines open questions in modern NLP Ignat et al. (2024b), and limitations in NLP and AI pipelines Mihalcea et al. (2025). Thus, our research goal in this work is threefold:
  1. RQ1:
    What NLP-based solutions already support positive social impact?
  2. RQ2:
    What are the main challenges in developing NLP for positive social impact?
  3. RQ3:
    What promising directions are currently overlooked?
To address these questions, we identify nine key NLP research directions, as illustrated in Figure 1, using the relevant SDGs and GRs. The following sections summarize research tasks, [^3] Details about methodology and motivation of paper selection can be found in the Appendix A. the open challenges and opportunities within each domain, concluding with overarching open research directions and recommended community actions for proactive NLP4SG efforts.
@@ 3. RQ3:
NLP in healthcare is a space that targets problems associated with SDG3: Good Health and Well-being by improving healthcare delivery and outcomes. It also targets critical global risks (GR4), including the decline in health and well-being, health workforce shortages, and infectious disease outbreaks.
Mental Health
Given that the need for mental health support increasingly remains unmet, NLP systems, specifically LLMs, address an access gap for users who are limited in time and resources Tong et al. (2023); Hua et al. (2024). As counselors, they could assist in detecting conditions like depression and addiction from clinical or social media data Giuntini et al. (2020); Yang et al. (2023a); responding empathetically by interpreting emotion and generating therapeutic dialogues Shen et al. (2020); Grandi et al. (2024); and tracking user mood or crises over time Čosić et al. (2024); Gong et al. (2019). As clients, NLP tools simulate diverse personas to train and evaluate counselors Louie et al. (2024); Liu et al. (2025a).
Physical Health
@@
Key challenges include health-related data, which is often scarce, sensitive, and affected by systemic biases, with language limitations and underrepresentation of marginalized groups raising ethical and privacy concerns (Ford et al., 2019; Shakeri Hossein Abad et al., 2022; Gunal et al., 2025); evaluation, which must go beyond accuracy to reflect fairness, contextual understanding, human-centered values (e.g. empathy), and ensure reproducibility; and long-term user impact, as reliance on LLMs for sensitive tasks risks sycophancy, ELIZA effects Ekbia (2008), and overdependence. There is also a lack of causal frameworks and interpretable models, which complicates the understanding in the outcomes of NLP systems Zhang et al. (2022). Remote care with LLM-based tools risks diverting individuals from essential in-person treatment Khawaja and Bélisle-Pipon (2023); Sweeney et al. (2021). This highlights a broader challenge: ensuring such technologies support—not replace—human professionals, with responsible use as a guiding principleShakeri Hossein Abad et al. (2022); Brown and Halpern (2021).
@@
Future work should extend beyond text to multimodal approaches that integrate speech prosody, facial expressions, physiological signals, and other data sources like sensors, wearables, or images for richer context and personalization Puce et al. (2025). Multi-agent and adaptive dialogue systems that consider user history, emotions, and culture could boost performance. Designing holistic evaluation frameworks in real-world simulations to ensure privacy, explainability, fairness, and accessibility can help address AI risks Lawrence et al. (2024); Yao et al. (2024). At the policy level, NLP can analyze public sentiment on AI used for healthcare applications from actual users, and health guidelines to inform better regulations and health campaigns Lindquist et al. (2021). Crucially, interdisciplinary research is needed to build AI-augmented therapeutic frameworks that complement human care—especially for underserved communities.
@@
The integration of LLMs into education offers tools that support SDG4: Quality Education, and also addresses a variety of risks by fostering informed and critical individuals (GR4, GR5). NLP is used for automated feedback Jurenka et al. (2024); Bauer et al. (2023); Gao et al. (2024); Stamper et al. (2024); Ramesh and Sanampudi (2022), tailored support Kazemitabaar et al. (2024); Daheim et al. (2024), and self-paced learning Kazemitabaar et al. (2023). NLP tools can expand access in underserved regions Yu et al. (2024), bridge language gaps Molina et al. (2024); Kwak and Pardos (2024), assist learners with disabilities Cheng et al. (2024), and ease teacher workload Lan and Chen (2024); Wang et al. (2024b); Shridhar et al. (2022).
AI Literacy
As AI systems become more embedded in everyday life, AI literacy is essential. A recent review (Yang et al., 2025) draws connections with digital, data, and algorithmic literacies. However, NLP-driven efforts for AI literacy remain scattered (Long and Magerko, 2020; Congress, 2023; Moorkens et al., 2024; Tapo et al., 2025; Korea Education and Research Information Service, 2025)
@@
Model limitations—such as lack of pedagogical reasoning Wang and Demszky (2023); Macina et al. (2023, 2025), misaligned explainability Okolo and Lin (2024), and accuracy Stamper et al. (2024); Kargupta et al. (2024)—hinder the effective integration of NLP into education. Mixed perceptions and mistrust toward AI also remain a barrier Nader et al. (2022); Laupichler et al. (2024). Broader issues like language, cultural differences, curriculum gaps, weak policy support, and limited infrastructure—especially in developing regions—further restrict equitable access to AI/NLP in education Kathala and Palakurthi (2025).
@@
Further research should be encouraged for specific groups, such as teachers Du et al. (2024), students Shen and Cui (2024), and other professionals Lo (2024). Another important opportunity is the alignment with expert-annotated pedagogical traces and grounding evaluations in curriculum-linked outcomes, to bridge the gap between subject expertise and pedagogical effectiveness Macina et al. (2025); Lucy et al. (2024). Human-in-the-loop approaches have also shown promise in scaling expert strategies while preserving teacher agency, particularly benefiting underserved communities Wang et al. (2024c). Multi-agent simulations offer privacy-preserving environments to test classroom policies and assess equity metrics before real-world deployment Zhang et al. (2025). For example, Socratic planning agents encourage critical thinking over rote memorization, fostering deeper learning experiences Kargupta et al. (2024). Finally, community-driven initiatives are crucial for consolidating best practices for the development of safe and accountable educational AI systems Chu et al. (2025); Wen et al. (2024).
@@
Economic downturn (GR3) is the sixth highest-ranked GRs, with serious implications for poverty—one of the world’s most pressing challenges (Lister, 2021) and the focus of SDG1. Nearly 700 million people continue to live on less than \$2.15 per day (Hasell et al., 2024; World Bank Group, 2024). NLP methods have been used to extract socioeconomic patterns from news (Lampos et al., 2014), analyze text (Paterson and Gregory, 2018; Hoeschle et al., 2025), classify poverty status Muñetón-Santa et al. (2022), extract poverty-related dimensions from interviews (Muñetón-Santa and Orozco-Arroyave, 2023) or analyze global narratives around poverty (Curto et al., 2024), and to identify performance disparities among different socioeconomic groups (Cercas Curry et al., 2024; Nwatu et al., 2023).
@@
To enable large-scale, global analysis of poverty, high-quality datasets labeled with income, socioeconomic status, or poverty indicators are essential, potentially gathered via data donations, user surveys, or social media statistics. [^4] https://datareportal.com/social-media-users Additionally, use-case-specific NLP applications can advance poverty research (Adauto et al., 2023). For example, model analysis can track the performance of current systems across socio-economic levels, information extraction can monitor government funding for poverty alleviation, and machine translation can improve resource access for non-English speakers. Finally, future work should develop clearer guidelines and taxonomies for categorizing poverty-related data and NLP methods to ensure consistency and comparability across studies.
@@
Peacebuilding is essential to achieving SDG16, and despite escalating threats like state-based armed conflict (R1); NLP’s application in key tools for peacebuilding like conflict prediction, human rights monitoring, and physical safety remains limited and underexplored.
Human Rights Violations
NLP in human rights mainly involves classification and detection of violations across languages, such as Arabic Twitter (Alhelbawy et al., 2016), Jordanian Arabic social media (Khalafat et al., 2021), Russian and Ukrainian Telegram (Nemkova et al., 2023), and English Twitter (Pilankar et al., 2022). Other tasks include threat detection for human rights defenders (Ran et al., 2023) and interpretable forced labor identification (Guzman et al., 2024). LLMs show potential for zero- and few-shot annotation but require human oversight (Nemkova et al., 2025b).
Conflict Prediction
NLP in conflict prediction mainly involves extracting and forecasting conflict dynamics using structured datasets like UCDP and ACLED (Hegre et al., 2019). Tasks include topic modeling for early signals (Mueller et al., 2024), modeling conflict stages with LLMs (Croicu and von der Maase, 2025; Nemkova et al., 2025a), automated report generation via retrieval-augmented generation (Nemkova et al., 2025c), and conflict content extraction and classification with domain-adapted models like ConfliBERT (Hu et al., 2022).
Physical Safety
NLP systems support understanding and intervention in threats to physical safety. They can be used to analyze social media, police reports, surveillance footage, health records, and news articles. Techniques like topic modeling More and Francis (2021), sentiment analysis Blevins et al. (2016), entity extraction Pavlick et al. (2016), and document classification Chang et al. (2018) help identify patterns of violence, perpetrators, and victims, informing interventions and policy decisions.
@@
First, Data and Domain Complexity: references to violations and threats are often indirect or euphemistic, complicating detection. Labeled multilingual datasets—especially for low-resource or sensitive regions—are scarce ALSaif and Alotaibi (2019); MacPhaul et al. (2023); Parker (2020); Chang et al. (2018), and causes are often multifaceted and extend beyond textual signals. Annotation is labor-intensive with expert disagreement common Blandfort et al. (2019); Levy et al. (2022); Botelle et al. (2022); Chew et al. (2023), while evaluation suffers from the absence of standardized benchmarks and reliable ground truth, limiting model validation and comparison. Second, Model Limitations: NLP models struggle with code-switching, generalization across languages, regions, and out-of-distribution data, and often reflect training biases. The black-box nature of LLMs reduces transparency, hindering their deployment in high-stakes contexts, with real-world application still in early stages and lacking validated best practices. Lastly, Ethical and Evaluation Challenges: misclassifications can lead to legal or advocacy risks, highlighting the need for rigorous evaluation and ethical care.
@@
On a computational level, tools such as conflict prediction models and systems for detecting human rights violations can support evidence-based policymaking, inform the development and targeting of interventions, and enable rapid programmatic adjustments in response to evolving conditions. Developing multilingual and cross-regional models will enable more comprehensive monitoring. Further, NLP methods can contribute to operational efficiency for implementing organizations. For example, LLMs offer capabilities for rapid document review and synthesis, facilitating timely and contextually informed analysis. RAG systems can further enhance information access by efficiently extracting relevant content from document repositories. Additionally, LLMs can support scenario generation, situational analysis, and strategic planning. Lastly, a dedicated peacebuilding workshop could offer a crucial forum for interdisciplinary collaboration on NLP tools for humanitarian aid and crisis response.
@@
NLP offers scalable tools for climate mitigation and adaptation—supporting key SDGs (SDG6, SDG11, SDG12, SDG14, SDG15) and addressing critical global environmental risks (R2)—by extracting insights from unstructured text like scientific papers, policy reports, and assessments (on Climate Change, 2022). Techniques such as topic modeling (Sietsma et al., 2023), summarization (Ghinassi et al., 2024), and classification (Varini et al., 2020; Stammbach et al., 2023; Bingler et al., 2022; Schimanski et al., 2023) enable analysis of diverse datasets. NLP can also help to detect misinformation and greenwashing by verifying claims (Diggelmann et al., 2020; Hsu et al., 2024). While fine-tuning or pretraining LLMs on climate text is common (Leippold et al., 2022; Thulke et al., 2024), RAG-based chatbots (Vaghefi et al., 2023) and automated fact-checkers (Leippold et al., 2025) are emerging rapidly.
@@
Extracting quantitative information, particularly from sustainability reports, remains challenging due to the complex, multi-modal and non-standardized nature of the data sources, leading to pipelines specifically designed for information extraction from tables (Mishra et al., 2024; Dimmelmeier et al., 2024). Furthermore, LLMs are prone to hallucination (Vaghefi et al., 2023) or are contaminated with false, conflicting or outdated climate information (Fore et al., 2024)Bulian et al. (2024) propose an evaluation framework based on presentational and epistemological qualities to evaluate quality and factual accuracy of LLMs.
@@
NLP is enabling cross-disciplinary collaboration, bridging domain experts with the fields of computer science, social science, and economics to analyze climate policies (Gandhi et al., 2024) or bringing the scientific community and non-governmental organization together to uncover narratives in public climate discourse (Gehring and Grigoletto, 2023; Rowlands et al., 2024). This is particularly important in expanding research to include cross-cultural and multilingual perspectives (Zhou et al., 2024; Bird et al., 2024).

### v2

# NLP for Social Good: A Survey and Outlook of Challenges, Opportunities, and Responsible Deployment

@@ Abstract
Natural language processing (NLP) now shapes many aspects of our world, yet its potential for positive social impact is underexplored. This paper surveys work in “NLP for Social Good" (NLP4SG) across nine domains relevant to global development and risk agendas, summarizing principal tasks and challenges. We analyze ACL Anthology trends, finding that inclusion and AI harms attract the most research, while domains such as poverty, peacebuilding, and environmental protection remain underexplored. Guided by our review, we outline opportunities for responsible and equitable NLP and conclude with a call for cross-disciplinary partnerships and human-centered approaches to ensure that future NLP technologies advance the public good.
NLP for Social Good: A Survey and Outlook of
@@
Antonia Karamolegkou1, Angana Borah2, Eunjung Cho3, Sagnik Ray Choudhury4, Martina Galletti5,6, Pranav Gupta7, Oana Ignat8, Priyanka Kargupta9, Neema Kotonya10, Hemank Lamba10, Sun-Joo Lee11, Arushi Mangla8, Ishani Mondal12, Fatima Zahra Moudakir13,14,15, Deniz Nazar16, Poli Nemkova4,11, Dina Pisarevskaya17, Naquee Rizwan18, Nazanin Sabri19, Keenan Samway13, Dominik Stammbach20, Anna Steinberg Schulten21,22, David Tomás23, Steven R Wilson24, Bowen Yi2,25, Jessica Zhu12, Arkaitz Zubiaga17, Anders Søgaard1, Alexander Fraser22,26, Zhijing Jin3,13,14,15, Rada Mihalcea2, Joel R. Tetreault10, Daryna Dementieva22,26 1University of Copenhagen 2University of Michigan-Ann Arbor 3ETH Zurich 4University of North Texas 5Sony Computer Science Laboratories - Paris 6University of Rome “La Sapienza” 7Lowe’s 8Santa Clara University 9University of Illinois Urbana-Champaign 10Dataminr 11United Nations Development Programme (UNDP) 12University of Maryland, College Park 13Max Planck Institute for Intelligent Systems, Tübingen 14Vector Institute 15University of Toronto 16University of Washington 17Queen Mary University of London 18IIT Kharagpur 19University of California San Diego 20Princeton University 21LMU Munich 22Munich Center for Machine Learning (MCML) 23University of Alicante 24University of Michigan-Flint 25University of Southern California 26Technical University of Munich Accepted to EACL 2026 (Main Conference), Rabat, Morocco. Correspondence: antoniakrm16@gmail.com, daryna.dementieva@tum.de
@@ 1 Introduction
To fully realize the potential of NLP, it is essential to look beyond technical achievements and reframe tasks around pressing societal needs. We draw on insights from the United Nations Sustainable Development Goals [^1] https://sdgs.un.org/goals (SDGs) and the 2025 Global Risks (GRs) Economic Report [^2] https://www.weforum.org/publications/global-risks-report-2025/digest to provide a foundation for an interdisciplinary recontextualization of NLP, encouraging reflection on how language technologies intersect with today’s most pressing challenges. We selected these two agendas as, from a social good perspective, UN SDGs offer a global framework for fostering peace and prosperity for people and the planet. However, while highly influential, these goals were established in 2015—prior to the rapid advancements in artificial intelligence. To contextualize them within today’s technological landscape, we also draw on insights from the 2025 GR Report, which highlights both the transformative potential and the emerging global risks associated with technology and information processing. The resulting mapping of NLP application domains to SDGs and GRs is shown in Figure 1. Our effort builds on prior research that assesses the role of NLP through positive impact Hovy and Spruit (2016); Jin et al. (2021), maps NLP4SG work to the SDGs Adauto et al. (2023); Gosselink et al. (2024), outlines open questions in modern NLP Ignat et al. (2024b), and limitations in NLP and AI pipelines Mihalcea et al. (2025). Thus, our research goal in this work is threefold: RQ1—what NLP-based solutions already support positive social impact, RQ2—what challenges arise in developing them, and RQ3—what promising directions remain overlooked?
We first analyze publication trends in the ACL Anthology across nine NLP research directions (Figure 2), revealing uneven growth across domains. For each domain, we review existing work and current challenges, [^3] Further statistics and details about survey methodology and paper selection are provided in Appendix B. and present an outlook on opportunities that reflects our perspective on how the field can advance. We conclude by synthesizing overarching research directions and community actions to encourage more proactive NLP4SG efforts.
Figure 2: Number of ACL Anthology papers per domain showing publication volume and growth trajectories over time (2019–2025).
@@
NLP can help address challenges linked to SDG3 (Good Health and Well-being) by improving healthcare delivery and outcomes; and to GR4 (Societal Risks), including health decline, health workforce shortages, and infectious disease outbreaks.
Mental Health.
As healthcare access remains uneven, NLP systems, particularly LLMs, offer scalable means to reduce support gaps for those limited in time and resources Tong et al. (2023); Hua et al. (2024). As counselors, they could assist in detecting conditions like depression and addiction from clinical or social media data Giuntini et al. (2020); Yang et al. (2023a); responding empathetically by interpreting emotion and generating therapeutic dialogues Shen et al. (2020); Grandi et al. (2024); and tracking user mood or crises over time Čosić et al. (2024); Gong et al. (2019). As clients, NLP tools simulate diverse personas to train and evaluate counselors Louie et al. (2024); Liu et al. (2025b).
Physical Health.
@@
Key challenges involve the availability of health-related data, which is scarce, sensitive, and often biased, with limited language coverage and marginalized group representation raising ethical and privacy concerns (Ford et al., 2019; Shakeri Hossein Abad et al., 2022; Gunal et al., 2025); the evaluation framework, which must go beyond accuracy to reflect fairness, contextual understanding, human-centered values (e.g. empathy), and ensure reproducibility; and long-term user impact, as reliance on LLMs for sensitive tasks risks sycophancy, ELIZA effects Ekbia (2008), and overdependence. There is also a lack of causal frameworks and interpretable models, which complicates the understanding in the outcomes of NLP systems Zhang et al. (2022). Remote care with LLM-based tools risks diverting individuals from essential in-person treatment Khawaja and Bélisle-Pipon (2023); Sweeney et al. (2021). This highlights a broader challenge: ensuring such technologies support—not replace—human professionals, with responsible use as a guiding principleShakeri Hossein Abad et al. (2022); Brown and Halpern (2021). There is growing evidence that LLMs lack the stability, contextual grounding, and ethical guarantees required for autonomous clinical or therapeutic decision-making (Zhao et al., 2024; Iftikhar et al., 2025; Cui et al., 2025), highlighting the need for informed, cautious, and supervised use of such tools.
@@
Future work should continue to emphasize multimodal approaches that jointly leverage text, speech prosody, facial expressions, physiological signals, and other data sources like sensors, wearables, or images for richer context and personalization Puce et al. (2025). Multi-agent and adaptive dialogue systems that consider user history, emotions, and culture could boost performance. Designing holistic evaluation frameworks in real-world simulations to ensure privacy, explainability, fairness, and accessibility can help address AI risks Lawrence et al. (2024); Yao et al. (2024). At the policy level, NLP can analyze public sentiment on AI used for healthcare applications from actual users, and health guidelines to inform better regulations and health campaigns Lindquist et al. (2021). Importantly, interdisciplinary research is needed to build AI-augmented therapeutic frameworks that complement human care—especially for underserved communities.
@@
The integration of LLMs into education offers tools that support SDG4: Quality Education, and also addresses a variety of societal risks by fostering informed and critical individuals (GR4). NLP is used for automated feedback Jurenka et al. (2024); Bauer et al. (2023); Gao et al. (2024); Stamper et al. (2024); Ramesh and Sanampudi (2022), tailored support Kazemitabaar et al. (2024); Daheim et al. (2024), and self-paced learning Kazemitabaar et al. (2023). NLP tools can expand access in underserved regions Yu et al. (2024), bridge language gaps Molina et al. (2024); Kwak and Pardos (2024), assist learners with disabilities Cheng et al. (2024), and ease teacher workload Lan and Chen (2024); Wang et al. (2024b); Shridhar et al. (2022). As AI systems become more embedded in everyday life, AI literacy is essential. A recent review (Yang et al., 2025) draws connections with digital, data, and algorithmic literacies. However, NLP-driven efforts for AI literacy remain scattered (Long and Magerko, 2020; Congress, 2023; Moorkens et al., 2024; Tapo et al., 2025; Korea Education and Research Information Service, 2025).
@@
Model limitations—such as lack of pedagogical reasoning Wang and Demszky (2023); Macina et al. (2023, 2025), misaligned explainability Okolo and Lin (2024), and accuracy Stamper et al. (2024); Kargupta et al. (2024) hinder the effective integration of NLP into education. Mixed perceptions and mistrust toward AI also remain a barrier Nader et al. (2022); Laupichler et al. (2024). Broader issues like language, cultural differences, curriculum gaps, weak policy support, and limited infrastructure—especially in developing regions—further restrict equitable access to AI/NLP in education Kathala and Palakurthi (2025).
@@
We believe LLMs can be a useful educational tool under sustained human oversight, but it cannot replace the relational, pedagogical, and contextual roles of human educators (Martynova et al., 2025; Calvert et al., 2025). Future work should focus on specific groups such as teachers Du et al. (2024), students Shen and Cui (2024), and other professionals Lo (2024) to obtain different perspectives on education. Moreover, aligning with expert-annotated pedagogical traces and grounding evaluations in curriculum outcomes to bridge subject expertise and pedagogical effectiveness Macina et al. (2025); Lucy et al. (2024). Human-in-the-loop methods can scale expert strategies while preserving teacher agency, particularly in underserved settings Wang et al. (2024c). Multi-agent simulations provide privacy-preserving environments to test classroom policies and equity before real-world deployment Zhang et al. (2025), for instance, through Socratic planning agents that promote critical thinking over rote learning Kargupta et al. (2024). Finally, community-driven efforts are key to consolidate best practices for safe and accountable educational AI Chu et al. (2025); Wen et al. (2024).
@@
Economic downturn (GR3) is the sixth highest-ranked GRs, with serious implications for poverty (SDG1)—one of the world’s most pressing challenges (Lister, 2021). Nearly 700 million people continue to live on less than \$2.15 per day (Hasell et al., 2024; World Bank Group, 2024). NLP methods have been used to extract socioeconomic patterns from news (Lampos et al., 2014), analyze text (Paterson and Gregory, 2018; Hoeschle et al., 2025), classify poverty status Muñetón-Santa et al. (2022), extract poverty-related dimensions from interviews (Muñetón-Santa and Orozco-Arroyave, 2023) or analyze global narratives around poverty (Curto et al., 2024), and to identify performance disparities among different socioeconomic groups (Cercas Curry et al., 2024; Nwatu et al., 2023).
@@
To enable a global analysis of poverty, datasets labeled with income, socioeconomic status, or poverty indicators are essential, potentially gathered via data donations, user surveys, or social media statistics. [^4] https://datareportal.com/social-media-users Additionally, use-case-specific NLP applications can advance poverty research (Adauto et al., 2023). For example, model analysis can track the performance of current systems across socio-economic levels, information extraction can monitor government funding for poverty alleviation, and machine translation can improve resource access for non-English speakers. Finally, future work should develop clearer guidelines and taxonomies for categorizing poverty-related data and NLP methods to ensure consistency and comparability across studies.
@@
Peacebuilding is essential to achieving SDG16, and despite escalating geopolitical threats (GR1), NLP tools for peacebuilding like human rights monitoring, conflict prediction, and physical safety remains limited and underexplored.
Human Rights Violations.
NLP for human rights focuses on detecting violations across languages and platforms, from Arabic and Jordanian social media (Alhelbawy et al., 2016; Khalafat et al., 2021) to Russian and Ukrainian Telegram (Nemkova et al., 2023) and English Twitter (Pilankar et al., 2022). Beyond general classification, efforts target threats to defenders (Ran et al., 2023a), forced labor with interpretable models (Guzman et al., 2024b), human trafficking (Liu et al., 2023; Saxena et al., 2023, 2024), dehumanization (Caporusso et al., 2024; Burovova and Romanyshyn, 2024), and other abuses (Yu, 2022; Guzman et al., 2024a; Sorato et al., 2024; De Paula et al., 2024; Wang, 2024).
Conflict Prediction.
NLP supports the extraction and forecasting of conflict dynamics (Halkia et al., 2020; Stoehr et al., 2021; Alsarra et al., 2023; Sathvik et al., 2024). Methods span from topic modeling for early warning (Mueller and Rauh, 2018; Mueller et al., 2024a, b) to LLM-based approaches modeling escalation and resolution, including emerging agentic paradigms (Croicu and von der Maase, 2025; Nemkova et al., 2025a; Nemkova and Albert, ). Studies also increasingly analyze civil unrest and protest dynamics (Sech et al., 2020; Chinta et al., 2021; Scharf et al., 2021; Raj et al., 2022; Siskou et al., 2022; Wiedemann et al., 2022; Loerakker et al., 2024; Olsen et al., 2024). Advances include domain-adapted encoders like ConfliBERT for improved extraction and classification (Hu et al., 2022) and automated situation reports via retrieval-augmentation (RAG) (Nemkova et al., 2025c). Recent work extends to diplomatic negotiation and argument mining, analyzing discourse in peace talks and diplomacy (Glaser et al., 2022; Rødven-Eide et al., 2023; Zaczynska and Stede, 2024; Zaczynska et al., 2024; Anisimova and Zikánová, 2024; Poiaganova and Stede, 2025) by modeling pathways to resolution.
Physical Safety.
They have been used to analyze relevant social media Al-Garadi et al. (2022a); Levy et al. (2022); Blandfort et al. (2019); Blevins et al. (2016), police reports Karystianis et al. (2019), surveillance footage Kumari et al. (2023), health records Borger et al. (2022); Botelle et al. (2022); MacPhaul et al. (2023), reporting systems Chew et al. (2023); Arseniev-Koehler et al. (2022), and news articles Pavlick et al. (2016). Techniques such as topic modeling More and Francis (2021), sentiment analysis Blevins et al. (2016), entity extraction Pavlick et al. (2016), and document classification Chang et al. (2018) help identify patterns of violence, perpetrators, and victims, informing interventions and policy decisions.
@@
Peacebuilding poses several challenges. Language and Context: threats and violations are often euphemistic, coded, or strategically disguised, especially in authoritarian or conflict settings Nemkova et al. (2023); Ran et al. (2023b); Nemkova et al. (2025b). Data is fragmented across low-resource languages and dialects, with annotation hindered by political sensitivity and expert disagreement Blandfort et al. (2019); Levy et al. (2022); Botelle et al. (2022); Chew et al. (2023). Temporal Volatility: conflicts evolve rapidly, as protests, ceasefires, or escalations outpace model adaptation, leaving systems trained on historical data prone to domain drift ALSaif and Alotaibi (2019); MacPhaul et al. (2023); Parker (2020); Chang et al. (2018). High-Stakes Ethics: misclassifications may expose activists, mislead humanitarian response, or legitimize violence, demanding stricter evaluation and oversight than other domains Kumari et al. (2023); Al-Garadi et al. (2022a); Chang et al. (2018). Since evidence for robust generalization remains limited, cautious interpretation and sustained human expertise are essential in deployment.
@@
We believe that models such as conflict predictors and human rights violation detectors can support policymaking, interventions, and programmatic adjustments across multilingual and cross-regional contexts ALSaif and Alotaibi (2019). NLP methods can also enhance operational efficiency: LLMs enable rapid document review and synthesis for timely, context-aware decisions Borger et al. (2022); Chew et al. (2023), while RAG systems improve access to relevant information Nemkova et al. (2025c). LLMs further assist with scenario generation, situational analysis, and strategic planning Nemkova et al. (2025c). Finally, a dedicated peacebuilding NLP workshop could provide a crucial interdisciplinary forum for developing tools for humanitarian aid and crisis response.
@@
NLP offers scalable tools for climate mitigation and adaptation—supporting key SDGs (SDG6, SDG7, SDG11, SDG12, SDG14, SDG15) and addressing critical global environmental risks (R2)—by extracting insights from unstructured text like scientific papers, policy reports, and assessments (on Climate Change, 2022). Techniques such as topic modeling (Sietsma et al., 2023), summarization (Ghinassi et al., 2024), and classification (Varini et al., 2020; Stammbach et al., 2023; Bingler et al., 2022; Schimanski et al., 2023) enable analysis of diverse datasets. NLP can also help to detect misinformation and greenwashing by verifying claims (Diggelmann et al., 2020; Hsu et al., 2024). While fine-tuning or pretraining LLMs on climate text is common (Leippold et al., 2022; Thulke et al., 2024), RAG-based chatbots (Vaghefi et al., 2023) and automated fact-checkers (Leippold et al., 2025) are emerging rapidly.
@@
Extracting quantitative information, particularly from sustainability reports, remains challenging due to the complex, multi-modal and non-standardized nature of the data sources, leading to pipelines specifically designed for information extraction from tables (Mishra et al., 2024; Dimmelmeier et al., 2024). Furthermore, LLMs, prone to hallucination (Vaghefi et al., 2023), often output false, conflicting or outdated climate information (Fore et al., 2024)Bulian et al. (2024) evaluate the quality and factual accuracy of LLMs on climate information using a framework based on presentational and epistemological qualities.
@@
NLP can enable cross-disciplinary collaboration, bridging domain experts from fields such as computer science, social science, and economics to analyze climate policies (Gandhi et al., 2024) or bringing the scientific community and non-governmental organizations to uncover narratives in public climate discourse (Gehring and Grigoletto, 2023; Rowlands et al., 2024). This is particularly important in expanding research to include cross-cultural and multilingual perspectives (Zhou et al., 2024; Bird et al., 2024). Lastly, NLP can inform other fields through climate reporting (Hershcovich et al., 2022b) and support sustainable behavior change (Chockkalingam et al., 2025).